OpenStudy (anonymous):

The mass of a proton is 1.00728 atomic mass units (amu) and the mass of a neutron is 1.00867 amu. What is the mass defect (in amu) of a nucleus whose nuclear mass is 59.9338 amu? What is the mass defect in kilograms? What is the energy equivalent of this mass in kilojoules? Confused on how to start. I know i multiply m by c^2.

OpenStudy (anonymous):

So would I add the mass of the proton and neutron onto the nuclear mass - then do the formula to find the total mass in Kilojoules?

OpenStudy (anonymous):

59.94975+1.00867+1.00729 = 61.94975 61.94975*(3*10^8)^2 = 5.575 *10^18? That would be the mass in KJ? maybe? lol

OpenStudy (chmvijay):

that's not mass defect

OpenStudy (chmvijay):

they haven't mentioned about which element hey are dealing with PLZ check once again ?

OpenStudy (chmvijay):

*they

OpenStudy (anonymous):

the mass nucleus of 60/27 Co nucleus - sorry not sure how that got left out!

OpenStudy (anonymous):

60 Co 27

OpenStudy (chmvijay):

LOL where di u leave this

OpenStudy (chmvijay):

*did

OpenStudy (anonymous):

The mass of a proton is 1.00728 atomic mass units (amu) and the mass of a neutron is 1.00867 amu. What is the mass defect (in amu) of a 60/27 Co nucleus whose nuclear mass is 59.9338 amu? What is the mass defect in kilograms? What is the energy equivalent of this mass in kilojoules?

OpenStudy (anonymous):

OOps added it back in!

OpenStudy (chmvijay):

ok now in 60 Co 27 can u tell me the no of protons and neutrons ?

OpenStudy (anonymous):

How would I figure that out?

OpenStudy (chmvijay):

no of protons = no of electrons = atomic no of element

OpenStudy (chmvijay):

LOL u have forgotten the fundamental of chem or what ?

OpenStudy (anonymous):

33 neutrons and 27 protons?

OpenStudy (anonymous):

The sum of both of those is 60.48247 right? By multypign each by their atomic mass?

OpenStudy (chmvijay):

yaaa

OpenStudy (chmvijay):

60.4847- 59.9338 = mass defect :) in amu unit

OpenStudy (anonymous):

.5509 - how do i convert to amu units?

OpenStudy (chmvijay):

now convert Amu unit to joules

OpenStudy (anonymous):

0.5509 to joules? So 0.5509 * (3*10^8)^2 = 4.958 *10^-16?

OpenStudy (chmvijay):

1MeV = 10^6 eV = 1.6 *10^(-13) J but 1 amu =931.5 MEV

OpenStudy (anonymous):

OH so I use 931.5 as the mass? I'm a bit confused

OpenStudy (chmvijay):

0.5509*931.5* 1.6 *10^(-13) gives ur answer in J LOL

OpenStudy (anonymous):

So the final answer would be 8.21*10^-11 joules of energy?

OpenStudy (anonymous):

and the mass defect would be 931.5? Or are the final answer and the mass defect the same thing? Sorry just a bit confused!

OpenStudy (anonymous):

@chmvijay whenever ya get time could you double check me up there ^^ thanks :)

OpenStudy (chmvijay):

how come ? 8.21*10^-11

OpenStudy (anonymous):

Is that the complete answer for the question? I'm just confused! 8.21*10^-11

OpenStudy (anonymous):

Thanks for all of your help!